40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Mar 11 15:19:45 2014, Jenne, Update, Electronics, High gain Trans PD electronics change IMG_1309.JPG
    Reply  Tue Mar 11 19:07:24 2014, ericq, Update, Electronics, High gain Trans PD electronics change SQRTINVwhitening.pdfsqrtinvWhitening.zip
    Reply  Thu Apr 17 16:04:40 2014, Jenne, Update, Electronics, High gain Trans PD electronics change 
Message ID: 9716     Entry time: Tue Mar 11 15:19:45 2014     Reply to this: 9720   9823
Author: Jenne 
Type: Update 
Category: Electronics 
Subject: High gain Trans PD electronics change 

As part of our CESAR testing last night, we had a look at the noise of the 1/sqrt(TR) signal. 

Looking at the time series data, while we were slowly sweeping through IR resonance (using the ALS), Rana noted that the linear range of the 1/sqrt(TR) signal was not as wide as it should be, and that this is likely because our SNR is really poor. 

When a single arm is at a normalized transmission power of 1, we are getting about 300 ADC counts.  We want this to be more like 3000 ADC counts, to be taking advantage of the full range of the ADC.

This means that we want to increase our analog gain by a factor of 10 for the low gain Thorlabs PDs. 

Looking at the photos from November when I pulled out the Xend transmission whitening board (elog 9367), we want to change "Rgain" of the AD620 on the daughter board.  While we're at it, we should also change the noisy black thick film resistors to the green thin film resistors in the signal path. 

The daughter board is D04060, S/N 101.  The main whitening board for the low gain trans QPD is D990399, RevB, S/N 104.

We should also check whether we're saturating somewhere in the whitening board by putting in a function generator signal via BNC cable into the input of the Thorlabs whitening path, and seeing where (in Dataviewer) we start to see saturation.  Is it the full 32,000 counts, or somewhere lower, like 28,000? 


Actually the gain was changed. From gain of 2 (Rgain = 49.4kOhm) to 20 (Rgain = 2.10kOhm), Corresponding calibration in CDS was also changed by locking the Xarm, running ASS, then setting the average arm power to be 1. Confirmed Xarm is locking. And now the signal is used for CESAR.  We see emperically that the noise has improved by a factor of approximately 10ish.

Attachment 1: IMG_1309.JPG  896 kB  Uploaded Tue Mar 11 23:08:18 2014  | Hide | Hide all
IMG_1309.JPG
ELOG V3.1.3-